OpenStudy (ash123456):

A rain gutter is to be constructed of aluminum sheets 12 inches wide. After marking off a length of 4 inches from each edge, this length is bent up at an angle θ. The area A of the opening may be expressed as the function: A(θ) = 16 sin θ ⋅ (cos θ + 1). If θ = 45°, what is the area of the opening?

OpenStudy (ash123456):

The answer choices are 4.8 in^2 9.0 in^2 19.3 in^2 20.8 in^2 I think it is 19.4 in^2

OpenStudy (ash123456):

*19.3

OpenStudy (sshayer):

correct,but can you show me your work?

OpenStudy (ash123456):

I converted the 45 degrees to radians using 45 times pi over 180. This got me 16 sin pi over 4 (cos pi over 4 + 1). Then, pi over 4 is 1 over the square root of 2. If you plug that in you get 16(1/sqrt2(1\sqrt2+1) and this simplifies to 8(1 + sqrt2) which is 19.3.

OpenStudy (sideswipe2085):

(10.06 MC) Functions f(x) and g(x) are shown below: f(x) g(x) f(x) = 3x2 + 12x + 16 g(x)=2sin(x-pie) Using complete sentences, explain how to find the minimum value for each function and determine which function has the smallest minimum y-value

OpenStudy (sideswipe2085):

@jfernandes

OpenStudy (sideswipe2085):

PLZZZZZ someone help me

OpenStudy (sshayer):

minimum value sin theta =-1 minimum value of g(x)=-2

OpenStudy (sshayer):

f(x)=3x^2+12x+16 =3(x^2+4x)+16 =3(x^2+4x+4-4)+16 =3(x^2+4x+4)-12+16 =3(x+2)^2+4 minimum value =?